OpenStudy (anonymous):

1/3y=7/3x+5/3 x-3y+5

OpenStudy (anonymous):

what exactly do you need help with?

OpenStudy (anonymous):

Solving the problem... @JM98SMITH

OpenStudy (phi):

is the first equation \[ \frac{1}{3} y = \frac{7}{3}x +\frac{5}{3} \]? and the 2nd "equation" is missing an = sign